OpenStudy (anonymous):

How do I verify that each given function is a solution of the given partial differential equation? a^2u_(xx)=u_t; u=(pi/t)^(1/2)e^(-x^2/4a^2t), t>0

OpenStudy (anonymous):

I know the function is kinda hard to see

myininaya (myininaya):

do you know hot to find u_x?

myininaya (myininaya):

\[u=(\frac{\pi}{t})^\frac{1}{2}e^{\frac{-x^2}{5a^2t}}\]

OpenStudy (anonymous):

No I don't dont really understand what u_xx is

myininaya (myininaya):

\[u=(\frac{\pi}{t})^\frac{1}{2}e^{\frac{-x^2}{4a^2t}}\] \[u_x=(\frac{\pi}{t})^\frac{1}{2} (\frac{-2x}{4a^2t})e^\frac{-x^2}{4a^2t}\]

myininaya (myininaya):

so i have found the partial derivative of u with respect to x

myininaya (myininaya):

now i just find the partial derivative of u_x with respect to x

OpenStudy (anonymous):

How do you find the partial derivative again?

myininaya (myininaya):

\[u_{xx}=(\frac{\pi}{t})^\frac{1}{2}\frac{-2}{4a^2t}e^{\frac{-x^2}{4a^2t}}+(\frac{\pi}{t})^\frac{1}{2}\frac{-2x}{4a^2t}(\frac{-2x}{4a^2t})e^{\frac{-x^2}{4a^2t}}\]

myininaya (myininaya):

you treat everything else like a constant except what you are taking the partial derivative with respect to

OpenStudy (anonymous):

so for the first one you just look at x and u since its u with respect to x?

myininaya (myininaya):

ok so now i'm going to rewrite u alittle so it is easier for me to take derivative of t (remember we are going to treat everything else like a constant except t since we are doing u_t) \[u=\pi^\frac{1}{2}t^{-\frac{1}{2}}e^\frac{-x^2t^{-1}}{4a^2}\] ....

OpenStudy (anonymous):

so wouldn't it be like taking the derivative of e^(1/t)?

myininaya (myininaya):

\[u_t=\pi^\frac{1}{2}\frac{-1}{2}t^{\frac{-3}{2}}e^\frac{-x^2t^{-1}}{4a^2}+\pi^\frac{1}{2}t^\frac{-1}{2}\frac{x^2t^{-2}}{4a^2}e^\frac{-x^2t^{-1}}{4a^2}\]

OpenStudy (anonymous):

Okay I'll try it first it does seem to make sense. Thanks for your help

OpenStudy (anonymous):

I'm confuse for the first one because wouldn't you need to do the product rule? because it's like (c/t)^(1/2) *(e^(c^-1))

OpenStudy (anonymous):

if you treat everything but t as a constant

myininaya (myininaya):

i used the product rule when i was doing the partial to u respect to t i also used the product rule when i was doing the partial to u_x with respect to x

OpenStudy (anonymous):

but for u with respect to t don't you get ____ + _____? like two terms?

myininaya (myininaya):

yes just like i did above

OpenStudy (anonymous):

?? I only see one term for u_x

OpenStudy (anonymous):

sorry if I'm confusing you

myininaya (myininaya):

you said the parital to t?

myininaya (myininaya):

so you are looking at the parital to x?

OpenStudy (anonymous):

never mind I just did the derivative for the first one wrong, did you use like substitution? I haven't done it in awhile

myininaya (myininaya):

yes once you find u_xx and u_t then plug into a^2u_xx=u_t

OpenStudy (anonymous):

Okay I'm just really confused with the partial derivative of u with respect to x

myininaya (myininaya):

this is everything i did:

OpenStudy (anonymous):

you got the derivative of u considering only t and everything is a constant right

myininaya (myininaya):

how do you find the derivative of \[f(x)=5e^{-\frac{x^2}{3}}\] ?

myininaya (myininaya):

u_x means you are treating everything else like a constant except x

OpenStudy (anonymous):

Oh I thought you meant t earlier

myininaya (myininaya):

\[f'(x)=5(\frac{-2x}{3})e^{\frac{-x^2}{3}}\]

myininaya (myininaya):

yes we found u_x and u_t earlier

OpenStudy (anonymous):

Oh I get it so for u_t you only look at t and for u_x you look at x

myininaya (myininaya):

yes for u_t you treat everything like it is a constant except t

OpenStudy (anonymous):

ohhh okay I get it thanks a lot for your help!! really appreciate it

myininaya (myininaya):

and for u_x you treat everything like a constant except x